# Gravitational lensing and shadows in the toron solution of Einstein’s equations using ray tracing methods

### v1

@@ 1. Introduction
The physical discussion of relativistic spacetimes in comparison to Newtonian gravity is best based on the tracing of light rays that are null geodesics in a relativistic setting but straight lines in a Newton theory. As already Einstein discussed in [10], light passing near strong gravitational sources on its way to an observer is subject to gravitational lensing which can lead to an apparent displacement of an object on the celestial sphere, the appearance of several copies of the same object and to an object directly behind the light bending source appearing as a ring-like structure called Einstein ring. It is also possible that no light from certain regions of the spacetime reaches an observer which leads to a shadow, see Synge [30] for the Schwarzschild and Bardeen [2] for the Kerr black hole. This can be due to photons actually orbiting the black hole along trajectories (which are circular orbits in the case of Schwarzshild, but can have a complicated shape in the case of Kerr [31]) restricted to a shell of spherical topology called the photon sphere (see the recent reviews [27, 24]).
@@ 2.1. Ernst equation
(3) $\mathcal{E}_{\xi\bar{\xi}}-\frac{1}{2(\bar{\xi}-\xi)}(\mathcal{E}_{\bar{\xi}}-\mathcal{E}_{\xi})=\frac{2}{\mathcal{E}+\bar{\mathcal{E}}}\mathcal{E}_{\xi}\mathcal{E}_{\bar{\xi}}\;.$
The metric functions (1) can be obtained from a given Ernst potential $\mathcal{E}$ via the formula $f=\Re\mathcal{E}$ and quadratures
(4) $A_{\xi}=2\rho\frac{(\mathcal{E}-\bar{\mathcal{E}})_{\xi}}{(\mathcal{E}+\bar{\mathcal{E}})^{2}},$
@@
(5) $k_{\xi}=(\xi-\bar{\xi})\frac{\mathcal{E}_{\xi}\bar{\mathcal{E}}_{\xi}}{(\mathcal{E}+\bar{\mathcal{E}})^{2}}\;.$
@@ 2.2. Kerr spacetimes
(6) $\mathcal{E}=\frac{\cos\varphi\,X-\mathrm{i}\sin\varphi\,Y-1}{\cos\varphi\,X-\mathrm{i}\sin\varphi\,Y+1},$
@@
The metric coefficients of the Kerr metric which corresponds to the Ernst potential (6) read
(7) $\begin{split}f&=\frac{\cos^{2}\varphi X^{2}+\sin^{2}\varphi Y^{2}-1}{(\cos\varphi X+1)^{2}+\sin^{2}Y^{2}},\\ e^{2k}&=\frac{m^{2}}{r_{+}r_{-}}\left(\sin^{2}\varphi(Y^{2}-1)+\cos^{2}\varphi(X^{2}-1)\right),\\ A&=\frac{2m\sin\varphi(1-Y^{2})(1+X\cos\varphi)}{\cos^{2}\varphi X^{2}+\sin^{2}\varphi Y^{2}-1}.\end{split}$
The metric is equatorially symmetric which means that all functions (7) are even in $z$.
@@
(8) $\mathrm{e}^{2k}=-\tan^{2}\varphi,\quad A=-1/\Omega_{BH},$
@@ 2.3. NUT spacetimes
(9) $\mathcal{E}=\frac{\tilde{X}-m-\mathrm{i}\ell}{\tilde{X}+m+\mathrm{i}\ell},$
@@
(10) $\begin{split}f&=\frac{\tilde{X}^{2}-(m^{2}+\ell^{2})}{(\tilde{X}+m)^{2}+\ell^{2}},\\ e^{2k}&=\frac{\tilde{X}^{2}-(m^{2}+\ell^{2})}{\tilde{X}^{2}-\tilde{Y}^{2}},\\ A&=\frac{2\ell}{\sqrt{m^{2}+\ell^{2}}}\tilde{Y}.\end{split}$
This solution can be interpreted as a black hole with real mass $m$ and gravimagnetic mass $\ell$. The latter is often refered to as the NUT parameter or gravimagnetic momentum in the literature, following the interpretation of [5] in which $\ell$ is interpreted as a massless source of angular momentum.
@@
Figure 2. Metric functions for the NUT solutions (10) for $m=\ell=1$, on the left $f$, on the right $e^{2k}$.
Similarly to the Kerr solution, there is an event horizon of spherical topology occupying the interval of the symmetry axis $z\in[-\sqrt{m^{2}+\ell^{2}},\sqrt{m^{2}+\ell^{2}}]$. The metric functions $f$ and $e^{2k}$ plotted in Fig. 2 are even in the $z$ variable, similarly to the case of the Kerr solution. This is no longer the case for the function $A$ given by (10) which is odd in $z$ (in contrast to the function $A$ for the Kerr solution given by (7) which is also even in $z$), see the left pane of Fig. 3. However, $A$ can be alternatively chosen to vanish on the positive (respectively negative) regular part of the axis by replacing $A$ by $A-2\ell$ (respectively $A+2\ell$), which would correspond to the choice of a rotating coordinate system. It is not possible in NUT spacetimes to choose $A$ to vanish on the whole regular part of the $z$-axis.
@@
Figure 3. Metric function $Af$ for the NUT solutions (10) for $m=\ell=1$, on the left for $A$ odd, on the right such that it vanishes for $z>\sqrt{m^{2}+\ell^{2}}$.
@@
which shows again that $A$ does not tend to 0 for $r\to\infty$ whereas this is the case for the Kerr solution. We get the following expression for the Ernst potential (9) on the axis:
@@ 3. Elliptic solutions to the stationary axisymmetric Einstein equations
The solutions from [20] are constructed as follows. Consider the family of elliptic curves $\mathcal{L}_{\xi}$ depending on spacetime variables $\rho$ and $z$ given by the equation
(11) $\mathcal{L}_{\xi}=\{(\lambda,\mu)\in\mathbb{C}^{2}\mid\mu^{2}=(\lambda-\mathrm{i}\sigma)(\lambda+\mathrm{i}\sigma)(\lambda-\xi)(\lambda-\bar{\xi})\},$
$\omega=\frac{1}{\mathcal{A}}\frac{d\lambda}{\mu},$
(12) $x(\lambda)=\frac{(\lambda+\mathrm{i}\sigma)(\mathrm{i}\sigma-\bar{\xi})}{(\lambda-\bar{\xi})2\mathrm{i}\sigma},\quad\lambda-\bar{\xi}=\frac{(\bar{\xi}+\mathrm{i}\sigma)(\mathrm{i}\sigma-\bar{\xi})}{2\mathrm{i}\sigma x-\mathrm{i}\sigma+\bar{\xi}},$
is real, positive and belongs to the interval $[1,\infty)$.
Then, introducing the elliptic module $k$ via
(13) $k^{2}=\frac{1}{x(\xi)}=\frac{4\rho\sigma}{z^{2}+(\rho+\sigma)^{2}},$
(which is real and $k\in[0,1]$) and the new variable $y$ via $x=y^{2}$ we get
Considering the complete elliptic integral of the first kind
$K(k)=\int_{0}^{1}\frac{dy}{[((y^{2}-1)(k^{2}y^{2}-1)]^{1/2}},$
we get for the period $\mathcal{A}$:
(14) $\mathcal{A}=\pm\frac{4K}{\sqrt{(\mathrm{i}\sigma-\bar{\xi})(\xi+\mathrm{i}\sigma)}}.$
The period of the curve $\tau=\int_{b}\omega$ reads
(15) $\tau=-\frac{1}{2\mathcal{A}\sqrt{(\mathrm{i}\sigma-\bar{\xi})(\xi+\mathrm{i}\sigma)}}\int_{0}^{\infty}\frac{dx}{\sqrt{x(1-x)(1-k^{2}x)}}=\frac{\mathrm{i}K^{\prime}}{K},$
where $K^{\prime}(k)=K(\sqrt{1-k^{2}})$.
For the integral of $\omega$ between $\xi$ and $\infty^{+}$ we have
(16) $\int_{\xi}^{\infty^{+}}\omega=-\frac{1}{2\mathcal{A}\sqrt{(\mathrm{i}\sigma-\bar{\xi})(\xi+\mathrm{i}\sigma)}}\int_{1/k^{2}}^{\frac{\mathrm{i}\sigma-\bar{\xi}}{2\mathrm{i}\sigma}}\frac{dx}{\sqrt{x(1-x)(1-k^{2}x)}}=\frac{1}{2K}(\gamma-K-\mathrm{i}K^{\prime})$
We need to compute the Abelian integral (16) for all values of $\rho>0$ and $z\in\mathbb{R}$. However, this integral is a multi-valued function of $\xi$ since it picks up a term $2\tau$ when $\xi$ goes around $\mathrm{i}\sigma$ (see Fig.2 of [21]). We are going to choose the discontinuity of the integral $\int_{\xi}^{\infty^{+}}\omega$ to appear for $0<\rho<\sigma$ in the equatorial plane.
The theta function with characteristic $\mathrm{p}\in\mathbb{R}$ and $\mathrm{q}\in\mathbb{C}$ is defined by
(17) $\vartheta_{\mathrm{pq}}(\mathrm{z},\tau)=\sum_{n\in\mathbb{Z}}\exp{(-\pi\mathrm{i}\tau(n+\mathrm{p})^{2}+2\pi\mathrm{i}(n+\mathrm{p})(\mathrm{z}+\mathrm{q}))},$
where $\mathrm{z}\in\mathbb{C}$ and $\tau\in\mathbb{H}$, i.e., $\tau$ is a complex number with $\Im(\tau)>0$. The theta function is usually defined as a function with a fixed period $\tau$, but in this paper both arguments depend on $(\xi,\bar{\xi})$ and therefore on the physical coordinates $\rho$ and $z$. Note that the theta function (17) with characteristics can be written in terms of the Jacobi theta function $\vartheta:=\vartheta_{3}=\vartheta_{00}$ with zero characteristics,
(18) $\vartheta_{\mathrm{pq}}(\mathrm{z},\tau)=e^{\mathrm{i}\pi{\tau}\mathrm{p}^{2}+2\pi\mathrm{i}\mathrm{p}(\mathrm{z}+\mathrm{q})}\vartheta(\mathrm{z}+2\pi\mathrm{i}({\tau}\mathrm{p}+\mathrm{q}),\tau)\;.$
The theta function can be computed in standard way (it is implemented in Matlab as well as other environments). Thus all of the above quantities can be computed in terms of standard elliptic functions, for instance in Matlab as used here.
With the above definitions, the elliptic solutions to the Ernst equation found in [19] can be written in the form
(19) $\mathcal{E}(\xi,\bar{\xi})=e^{-\pi\mathrm{i}\mathrm{p}}\frac{\vartheta_{\mathrm{pq}}(\int^{\infty^{+}}_{\xi}\omega,\tau)}{\vartheta_{\mathrm{pq}}(\int^{\infty^{-}}_{\xi}\omega,\tau)},$
where $\int^{\infty^{-}}_{\xi}\omega=-\int^{\infty^{+}}_{\xi}\omega$. The constants $\mathrm{p}$ and $\mathrm{q}$ satisfy the reality conditions $\mathrm{p}\in\mathbb{R}$ and $\mathrm{q}\in\mathrm{i}\mathbb{R}$  [^1] In the original paper [16] the necessary unitary constant factor $e^{-\pi\mathrm{i}\mathrm{p}}$ in (19) was omitted; this mistake was corrected in [9]..
Note that the dependence of the Ernst potential on the physical coordinates is via the branch points $\xi$ and $\bar{\xi}$ of the family of elliptic curves (11). Moreover, this family of solutions is parametrized by $\mathrm{p}\in\mathbb{R}$ and $\mathrm{q}\in\mathrm{i}\mathbb{R}$. For the numerical analysis of the solution one can use the approach of [13], but here we alternatively use the elliptic functions implementation in Matlab.
Without loss of generality, we will assume that $\sigma=1$, i.e., the fixed branch points are placed at $\mathrm{i}$ and $-\mathrm{i}$. A different choice of $\sigma\neq 0$ corresponds to a constant rescaling of $\xi$.
@@
The fact that the integral $\int^{\infty^{+}}_{\xi}\omega$ is branched in the interval $0\leq\rho\leq 1$ of the equatorial plane implies that the Ernst potential is discontinuous for $10<\rho<$ 1 as can be seen in Fig. 4. In contrast to the disk discussed in [17] the metric is not continuous at the disk which makes an interpretation as a delta-type distribution, for instance a dust disk, problematic since the energy-momentum tensor would be proportional to the derivative of a delta function in this case. One possible interpretation of this solution is thus that it is defined on an infinitely-sheeted spacetime whose slice for constant $\phi$ and $t$ is the Riemann surface in $\xi$-coordinate whose “branch ring” is located at $z=0$, $\rho=1$, see [20]. In this paper we shall be considering only the first “sheet” of the spacetime ignoring its non-trivial topology and, therefore, we will treat the disk $0\leq\rho\leq 1$ as in [8] as totally light absorbing. Thus we discuss the spacetime with a discontinuity at the interval $z=0$, $0\leq\rho\leq 1$.
@@
Figure 4. Ernst potential (19) for $\alpha=0.3$, $\mathrm{p}=0$. On the left the real part, on the right the imaginary part.
Below we also put $\mathrm{p}=0$ which simplifies the branching structure and the behavior of solution on the axis. We also denote
$\mathrm{q}=\mathrm{i}\alpha$
for $\alpha\in\mathbb{R}_{+}$. Then (19) coincides with the toron solution of [20]:
(20) $\mathcal{E}(\xi,\bar{\xi})=\frac{\vartheta(\int^{\infty^{+}}_{\xi}\omega+\mathrm{i}\alpha,\tau)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha,\tau)}\;.$
The metric functions corresponding to the Ernst potential (20) can be given in closed form, see [16]:
(21) $f=Q\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\mathrm{i}\alpha+1/2)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)},$
(22) $e^{2k}=\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\mathrm{i}\alpha+1/2)}{\vartheta(0)\vartheta(1/2)},$
(23) $(A-A_{0})f=-\rho\left(\frac{1}{Q}\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\xi}\omega+\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)}-1\right)$
@@
(24) $Q=\frac{\vartheta(\int^{\infty^{-}}_{\xi}\omega)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega)}{\vartheta(0)\vartheta(1/2)}.$
@@ 3.1. Physical properties
We are interested in providing a physical interpretation of the parameter $\alpha\in\mathbb{R}_{+}$, which is done by studying the asymptotic behaviour of the metric and by studying light rays in various spacetimes in a later section. Before that, we describe the peculiar ergospheres in these spacetimes.
Let us recall that the ergosphere is defined as the hypersurface where the metric component $g_{tt}=-f$ vanishes. From the explicit form of the real part of the Ernst potential and the form of the zeros of theta functions, one can see that $f$ can only vanish if the theta function $\vartheta(\mathrm{i}\alpha+1/2,\tau)$ vanishes. This condition is equivalent to the existence of $n,m\in\mathbb{N}$ such that
$1/2+\mathrm{i}\alpha=(m+1/2)+(n+1/2)\tau.$
Thus $m=0$, and for each period
(25) $\tau_{n}=\frac{\mathrm{i}\alpha}{n+1/2}\;,\hskip 19.91684ptn=0,1,\dots$
we get a separate component of the ergosphere. The set of coordinates $\rho,z$ corresponding to the period $\tau_{n}$ can be described via the equation (13) and by considering the inverse of (14), see [6], which is given by
$\kappa^{2}=\frac{\vartheta_{2}^{4}(0,\tau)}{\vartheta_{3}^{4}(0,\tau)},$
where $\vartheta_{2}$ and $\vartheta_{3}$ are Jacobi elliptic functions.
Denote $\kappa_{n}=\kappa(\tau_{n})$ for $n\in\mathbb{N}$. Then the equation (25) for the $n$th component of the ergosphere is equivalent to the equation (see (13))
$\kappa_{n}^{2}=\frac{4\rho\sigma}{z^{2}+(\rho+\sigma)^{2}}$
which leads to
@@
where
$\rho_{n}=\frac{2}{\kappa_{n}^{2}}-1\;,\hskip 19.91684ptR_{n}=2\frac{\sqrt{1-\kappa_{n}^{2}}}{\kappa_{n}^{2}}\;.$
Therefore, for each value of the parameter $\alpha$ (which parametrizes the family of elliptic solutions), there is an infinite family of ergospheres parametrized by $n\in\mathbb{N}$ and since the spacetime is stationary axisymmetric, the spatial projection of each ergosphere is a torus whose major radius is $\rho_{n}$ and whose minor radius is $R_{n}$.
@@
The size of every ergosphere increases in dependence of $\alpha$, i.e. both radii $\rho_{n}$ and $R_{n}$ increase monotonically with $\alpha$ for every $n\in\mathbb{N}$, which can be deduced from the explicit formulas shown above. This can be observed in Table 1 for $E_{0}$.
@@
On the axis ($\rho\to 0$) the branch cut $[\xi,\bar{\xi}]$ collapses to a point, and all quantities on the elliptic curves (11) can be expressed in terms of quantities defined on the genus 0 curve $\mu^{2}=\lambda^{2}+1$, see for instance [18] and references therein. For small $\rho$, the $b$-period has an expansion of the form $\tau=2\ln\rho+\sum_{n=0}^{\infty}c_{n}\rho^{n}$ with constants $c_{n}$, $n=0,1,\ldots$.
@@
To study the asymptotic behaviour for the elliptic Ernst potential, we are going to compute the potential on the axis when the branch points $\xi$ and $\bar{\xi}$ coincide and the elliptic curve degenerates.
(26) $\tau=-\pi\mathrm{i}\ln\rho+\dots$
Therefore,
@@
Moreover,
As the period of the curve tends to infinity according to (26) only two terms in the series (18) remain non-zero. Namely,
Therefore, we have
(27) $\mathcal{E}(z,\rho=0)=\frac{z+\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{-2\pi\alpha}}{z+\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{2\pi\alpha}}\;,\hskip 19.91684ptz\leq 0.$
Similarly, for $z\geq 0$ the differential $\omega$ has residue $-1/2\pi\mathrm{i}$ at $z^{+}$ and the residue $-1/2\pi\mathrm{i}$ at $z^{-}$ (Fig. 8).
Thus, as $\rho\to 0$,
and
Therefore,
and
(28) $\mathcal{E}(z,\rho=0)=\frac{z-\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{-2\pi\alpha}}{z-\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{2\pi\alpha}}\;,\hskip 19.91684ptz\geq 0.$
The expressions (27) and (28) imply
(29) $\mathcal{E}=1-\frac{\mathrm{i}\sigma}{z}\sinh(2\pi\alpha)+\mathcal{O}(1/z^{2})\;;$
therefore, the “mass” $\mathrm{i}\sigma\sinh(2\pi\alpha)$ is purely imaginary in the elliptic solution. In the sequel, we refer to this solution as the toron spacetime.

### v2

@@ 1. Introduction
The physical discussion of relativistic spacetimes in comparison to Newtonian gravity is best based on the tracing of light rays that are null geodesics in a relativistic setting but straight lines in a Newton theory. As already Einstein discussed in [10], light passing near strong gravitational sources on its way to an observer is subject to gravitational lensing which can lead to an apparent displacement of an object on the celestial sphere, the appearance of several copies of the same object and to an object directly behind the light bending source appearing as a ring-like structure called Einstein ring. It is also possible that no light from certain regions of the spacetime reaches an observer which leads to a shadow, see Synge [30] for the Schwarzschild and Bardeen [2] for the Kerr black hole. This can be due to photons actually orbiting the black hole along trajectories (which are circular orbits in the case of Schwarzschild, but can have a complicated shape in the case of Kerr [31]) restricted to a shell of spherical topology called the photon sphere (see the recent reviews [27, 24]).
@@ 2.1. Ernst equation
(2.3) $\mathcal{E}_{\xi\bar{\xi}}-\frac{1}{2(\bar{\xi}-\xi)}(\mathcal{E}_{\bar{\xi}}-\mathcal{E}_{\xi})=\frac{2}{\mathcal{E}+\bar{\mathcal{E}}}\mathcal{E}_{\xi}\mathcal{E}_{\bar{\xi}}\;.$
The metric functions (2.1) can be obtained from a given Ernst potential $\mathcal{E}$ via the formula $f=\Re\mathcal{E}$ and quadratures
(2.4) $A_{\xi}=2\rho\frac{(\mathcal{E}-\bar{\mathcal{E}})_{\xi}}{(\mathcal{E}+\bar{\mathcal{E}})^{2}},$
@@
(2.5) $k_{\xi}=(\xi-\bar{\xi})\frac{\mathcal{E}_{\xi}\bar{\mathcal{E}}_{\xi}}{(\mathcal{E}+\bar{\mathcal{E}})^{2}}\;.$
@@ 2.2. Kerr spacetimes
(2.6) $\mathcal{E}=\frac{\cos\varphi\,X-\mathrm{i}\sin\varphi\,Y-1}{\cos\varphi\,X-\mathrm{i}\sin\varphi\,Y+1},$
@@
The metric coefficients of the Kerr metric which corresponds to the Ernst potential (2.6) read
(2.7) $\begin{split}f&=\frac{\cos^{2}\varphi X^{2}+\sin^{2}\varphi Y^{2}-1}{(\cos\varphi X+1)^{2}+\sin^{2}Y^{2}},\\ e^{2k}&=\frac{m^{2}}{r_{+}r_{-}}\left(\sin^{2}\varphi(Y^{2}-1)+\cos^{2}\varphi(X^{2}-1)\right),\\ A&=\frac{2m\sin\varphi(1-Y^{2})(1+X\cos\varphi)}{\cos^{2}\varphi X^{2}+\sin^{2}\varphi Y^{2}-1}.\end{split}$
The metric is equatorially symmetric which means that all functions (2.7) are even in $z$.
@@
(2.8) $\mathrm{e}^{2k}=-\tan^{2}\varphi,\quad A=-1/\Omega_{BH},$
@@ 2.3. NUT spacetimes
(2.9) $\mathcal{E}=\frac{\tilde{X}-m-\mathrm{i}\ell}{\tilde{X}+m+\mathrm{i}\ell},$
@@
(2.10) $\begin{split}f&=\frac{\tilde{X}^{2}-(m^{2}+\ell^{2})}{(\tilde{X}+m)^{2}+\ell^{2}},\\ e^{2k}&=\frac{\tilde{X}^{2}-(m^{2}+\ell^{2})}{\tilde{X}^{2}-\tilde{Y}^{2}},\\ A&=\frac{2\ell}{\sqrt{m^{2}+\ell^{2}}}\tilde{Y}.\end{split}$
This solution can be interpreted as a black hole with real mass $m$ and gravimagnetic mass $\ell$. The latter is often referred to as the NUT parameter or gravimagnetic momentum in the literature, following the interpretation of [5] in which $\ell$ is interpreted as a massless source of angular momentum.
@@
Figure 2. Metric functions for the NUT solutions (2.10) for $m=\ell=1$, on the left $f$, on the right $e^{2k}$.
Similarly to the Kerr solution, there is an event horizon of spherical topology occupying the interval of the symmetry axis $z\in[-\sqrt{m^{2}+\ell^{2}},\sqrt{m^{2}+\ell^{2}}]$. The metric functions $f$ and $e^{2k}$ plotted in Fig. 2 are even in the $z$ variable, similarly to the case of the Kerr solution. This is no longer the case for the function $A$ given by (2.10) which is odd in $z$ (in contrast to the function $A$ for the Kerr solution given by (2.7) which is also even in $z$), see the left pane of Fig. 3. However, $A$ can be alternatively chosen to vanish on the positive (respectively negative) regular part of the axis by replacing $A$ by $A-2\ell$ (respectively $A+2\ell$), which would correspond to the choice of a rotating coordinate system. It is not possible in NUT spacetimes to choose $A$ to vanish on the whole regular part of the $z$-axis.
@@
Figure 3. Metric function $Af$ for the NUT solutions (2.10) for $m=\ell=1$, on the left for $A$ odd, on the right such that it vanishes for $z>\sqrt{m^{2}+\ell^{2}}$.
@@
which shows again that $A$ does not tend to 0 for $r\to\infty$ whereas this is the case for the Kerr solution. We get the following expression for the Ernst potential (2.9) on the axis:
@@ 3. Elliptic solutions to the stationary axisymmetric Einstein equations
The metric functions describing the toron spacetimes are given in terms of theta functions and Abel maps on elliptic curves. In this section we shall simply present the metric, the reader is referred to Appendix A for definitions. All elliptic functions can be computed in standard way (it is implemented in Matlab as well as other environments).
We remark that the dependence of the Ernst potential on the physical coordinates is via the branch points $\xi$ and $\bar{\xi}$ of the family of elliptic curves (A.1).
@@
It is important to note that the integral $\int^{\infty^{+}}_{\xi}\omega$ is branched in the interval $0\leq\rho\leq 1$ of the equatorial plane, which implies that the Ernst potential is discontinuous for $10<\rho<$ 1 as can be seen in Fig. 4. In contrast to the disk discussed in [17] the metric is not continuous at the disk which makes an interpretation as a delta-type distribution, for instance a dust disk, problematic since the energy-momentum tensor would be proportional to the derivative of a delta function in this case. One possible interpretation of this solution is thus that it is defined on an infinitely-sheeted spacetime whose slice for constant $\phi$ and $t$ is the Riemann surface in the $\xi$-coordinate whose “branch ring” is located at $z=0$, $\rho=1$, see [20]. In this paper we shall be considering only the first “sheet” of the spacetime ignoring its non-trivial topology and, therefore, we will treat the disk $0\leq\rho\leq 1$ as in [8] as totally light absorbing. Thus we discuss the spacetime with a discontinuity at the interval $z=0$, $0\leq\rho\leq 1$. The Ernst potential for the toron solution reads
(3.1) $\mathcal{E}(\xi,\bar{\xi})=\frac{\vartheta(\int^{\infty^{+}}_{\xi}\omega+\mathrm{i}\alpha,\tau)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha,\tau)}\;.$
We show an example in Fig. 4.
@@
Figure 4. Ernst potential (A.9) for $\alpha=0.3$, $\mathrm{p}=0$. On the left the real part, on the right the imaginary part.
The metric functions corresponding to the Ernst potential (3.1) can be given in closed form, see [16]:
(3.2) $f=Q\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\mathrm{i}\alpha+1/2)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)},$
(3.3) $e^{2k}=\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\mathrm{i}\alpha+1/2)}{\vartheta(0)\vartheta(1/2)},$
(3.4) $(A-A_{0})f=-\rho\left(\frac{1}{Q}\frac{\vartheta(\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\xi}\omega+\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)}{\vartheta(\int^{\infty^{-}}_{\xi}\omega+\mathrm{i}\alpha)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega+\mathrm{i}\alpha)}-1\right),$
@@
(3.5) $Q=\frac{\vartheta(\int^{\infty^{-}}_{\xi}\omega)\vartheta(\int^{\infty^{-}}_{\bar{\xi}}\omega)}{\vartheta(0)\vartheta(1/2)}.$
@@ 3.1. Physical properties
We are interested in providing a physical interpretation of the parameter $\alpha\in\mathbb{R}_{+}$, which is done by studying the asymptotic behavior of the metric and by studying light rays in various spacetimes in a later section. Before that, we describe the peculiar ergospheres in these spacetimes.
Let us recall that ergospheres are defined as the hypersurfaces satisfying the condition $g_{tt}=0$. Given a spacetime with $\alpha$, it is shown in Appendix A.1 that this spacetime admits the infinite family of ergospheres described by
@@
for all $n\in\mathbb{N}$, where $\rho_{n}$ and $R_{n}$ are functions that increase monotonically with the magnitude of $\tau_{n}=\mathrm{i}\alpha/(n+\frac{1}{2})$, meaning that their value in dependence of $n$ is decreasing. The explicit form of both radii is given by (A.11) in Appendix A.1.
Therefore, for each value of the parameter $\alpha$ (which parametrizes the class of elliptic solutions), there is an infinite family of ergospheres parametrized by $n\in\mathbb{N}$ and since the spacetime is stationary axisymmetric, the spatial projection of each of these surfaces is a torus with major radius $\rho_{n}$ and minor radius $R_{n}$. As mentioned above, both of these radii decrease monotonically in dependence of $n$; moreover, as $n$ tends to infinity we observe that $\rho_{n}\to 1$ and $R_{n}\to 0$, i.e., the limiting ergosphere is a ring of radius 1.
@@
The size of the main ergosphere $E_{0}$ increases in dependence of $\alpha$, i.e. both radii $\rho_{0}$ and $R_{0}$ increase monotonically with $\alpha$ (the same can be said about the $n$th ergosphere, for every $n\in\mathbb{N}$), which can be deduced from the explicit formulas presented in Appendix A.1. This monotonic increase can be observed in Table 1 for $E_{0}$ for some values of $\alpha$.
@@
On the axis ($\rho\to 0$) the branch cut $[\xi,\bar{\xi}]$ collapses to a point, and all quantities on the elliptic curves (A.1) can be expressed in terms of quantities defined on the genus 0 curve $\mu^{2}=\lambda^{2}+1$, see for instance [18] and references therein. For small $\rho$, the $b$-period has an expansion of the form $\tau=2\ln\rho+\sum_{n=0}^{\infty}c_{n}\rho^{n}$ with constants $c_{n}$, $n=0,1,\ldots$.
@@
To study the asymptotic behavior for the elliptic Ernst potential, this value is computed analytically on the axis when the branch points $\xi$ and $\bar{\xi}$ coincide and the elliptic curve degenerates. The Ernst potential on the axis (the step-by-step computation is presented in Appendix A.1) is described by the elementary functions
(3.6) $\mathcal{E}(z,\rho=0)=\frac{z+\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{-2\pi\alpha}}{z+\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{2\pi\alpha}}\;,\hskip 19.91684ptz<0,$
@@
(3.7) $\mathcal{E}(z,\rho=0)=\frac{z-\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{-2\pi\alpha}}{z-\sqrt{z^{2}+\sigma^{2}}-\mathrm{i}\sigma e^{2\pi\alpha}}\;,\hskip 19.91684ptz>0.$
The expressions (3.6) and (3.7) imply the expansion
(3.8) $\mathcal{E}=1-\frac{\mathrm{i}\sigma}{z}\sinh(2\pi\alpha)+\mathcal{O}(1/z^{2})\;$
around infinity. Therefore, the “mass” $\mathrm{i}\sigma\sinh(2\pi\alpha)$ is purely imaginary in the elliptic solution. In the sequel, we refer to this solution as the toron spacetime.